OpenStudy (anonymous):

The town of Clueville has approximately 30 000 people. An arts major is researching the spread of news across the town. She determines that 300 knew the news of Col. Mustard's death the morning after it happened. After 1 day, she found that 900 people knew about the death. Solve the logistic equation dy/dt = k*y( 1 - y/L),to estimate how many people in the town knew about the death after 4 days. Note: L is the total possible number of people who can know and k is the growth constant.

OpenStudy (anonymous):

i think you need to integrate in this question.

OpenStudy (anonymous):

every day the number of people who know triples

OpenStudy (anonymous):

so you can use \(300\times 3^t\) where \(t\) is the number of days after the first day. i have no idea what that other stuff is about, but whatever you get will be equivalent to \(300\times 3^t\)

OpenStudy (anonymous):

since you are interested in after four days, we need no formula at all day 0: 300 day 1: 900 day 2: 2700 day 3: 8100 day 4: whatever \(8100\times 3\) is

OpenStudy (anonymous):

So what do I do with 300x3^t

hartnn (hartnn):

u just calculate 3*8100, that would be your answer

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Its suppose to be a differential equation problem, I don't use DE's anywhere?

hartnn (hartnn):

oh, i just followed satellite.... no, then u have to use diff. equation

OpenStudy (anonymous):

How do I do it using differential equations. Can you take me through it?

hartnn (hartnn):

use variable separable method,this is easily separable bring dy and y terms on one side of = sign and dt term on other side, what u get ?

OpenStudy (anonymous):

dy/y(1 - y/2) = k*dt, is that right?

hartnn (hartnn):

why L =2 ? keep it as L. otherwise its correct !

OpenStudy (anonymous):

oh sorry!

OpenStudy (anonymous):

Then I integrate?

OpenStudy (anonymous):

What do I do with L?, its the total possible number of people that can know. What do I do with it in the integral

hartnn (hartnn):

L is constant. it won't be a trouble while integration. u need to integrate 1/y(1 - y/L) u know how ,right ?

OpenStudy (anonymous):

can you do it? cause I've been stuck on this forever. I'm making a mistake and I dont see it.

hartnn (hartnn):

ok, u know partial fractions ?

OpenStudy (anonymous):

not very well

hartnn (hartnn):

write that as A/y + B / (1-y/L) = 1/y(1 - y/L) find A and B

hartnn (hartnn):

so u get 1/y - 1/ (y-L)

hartnn (hartnn):

now integrate {1/y - 1/ (y-L)} dy

hartnn (hartnn):

i think u can, that is easy

OpenStudy (anonymous):

Sorry, I'm not getting it :/ I suck at partial fractions

OpenStudy (anonymous):

integral y-29999ln(y-1)?